Phathu needs R10 500 in ten months’ time to buy herself a new lens for her camera. Two months ago she deposited R9000 in a savings account at a simple interest rate of 11,5% per year. How much money will

 Phathu

still need to buy the lens ten months from now?


1
Expert's answer
2021-05-21T03:14:01-0400

Present money=Principal deposited+Interest amount

=9000+{(9000*11.5*2/12)}/100

=9000+172.5

=9172.5

Money required =10500-9172.50=R1327.5